Kroger Store decides to run a price promotion for 20oz bottle of Coca Cola. They know...

  1. Kroger Store decides to run a price promotion for 20oz bottle of Coca Cola. They know the elasticity of demand for Coke is -4. During the price promotion, they expect to sell 28% more bottles. Before the price drop, they sold 100,000 bottles.
    1. For them to be able to achieve this increases sale of bottles, what would the percentage change in price be?
  1. If the price before the promotion was $1.00 per bottle, what is the new price per bottle?
  1. Calculate the revenue before the price drop.
  1. Calculate the revenue after the price drop.
  1. Is the price drop a good marketing strategy, given all else is same?
Answer #1

a)

Percent change in price=Percent change in quantity demanded/Price elasticity of demand

Percent change in price=28%/(-4)=-7%

Price should decrease by 7%

b)

Given initial price=P1=$1

Promotional price=P2=1*(1-7%)=$0.93

c)

Revenue before price drop=P1*Q1=1*100000=$100,000

d)

Quantity after price drop=100000*(1+28%)=128,000

Revenue after price drop=P2*Q2=0.93*128000=$119040

e)

There is an increase in revenue. Its a good strategy.